OpenStudy (anonymous):

What is the function rule for the graph shown? A graph of the ordered pairs: 9 comma zero, eleven comma 2, twelve comma 3, and fourteen comma 5. y = 9x y = x + 9 y = x – 9 y equals one over nine times x

OpenStudy (anonymous):

@TylerD @SolomonZelman

OpenStudy (tylerd):

you should attempt to learn this stuff before taking the quiz

OpenStudy (anonymous):

thx for the advise

OpenStudy (tylerd):



OpenStudy (tylerd):

one sec

OpenStudy (tylerd):

so i believe its (x,y)

OpenStudy (tylerd):

for the first one, if x=9 y=0

OpenStudy (tylerd):

because this is multiple choice

OpenStudy (tylerd):

we actually dont need to know all the math, you can pretty much figure it out by guess and check, if you know the material to begin with.

OpenStudy (tylerd):

if y=0 and x=9 lets check the first one y=9x 0=9 times 9

OpenStudy (tylerd):

0 doesnt equal 9x so it cant be that one

OpenStudy (anonymous):

i think its B

OpenStudy (tylerd):

0=9+9

OpenStudy (tylerd):

nope

OpenStudy (anonymous):

NO sorry i meant

OpenStudy (anonymous):

C

OpenStudy (tylerd):

it works for 0=9-9

OpenStudy (tylerd):

lets see if it works for the others

OpenStudy (anonymous):

OK

OpenStudy (tylerd):

(11,2) 2=11-9

OpenStudy (tylerd):

yep

OpenStudy (tylerd):

but

OpenStudy (anonymous):